OpenStudy (anonymous):

In Triangle ABC, Angle B=90*, SinC=3/5. AB=9cm. Find BC.

jhonyy9 (jhonyy9):

can you drawing it ? this will help you sure

OpenStudy (anonymous):

@jhonyy9 - I have it on my paper...... should i draw it here?

jhonyy9 (jhonyy9):

yes please

jhonyy9 (jhonyy9):

just for i can explaning to you for you can understanding it easy

OpenStudy (anonymous):

 Okay? @jhonyy9

jhonyy9 (jhonyy9):

can you writing here what mean sinC ,sinus of angle C , using proportionality of sides ?

OpenStudy (anonymous):

SinC=AB/BC?

jhonyy9 (jhonyy9):

exactly so than now just you need rewriting it and let be equale by 3/5 and from this you njust need calculi BC ok ?

jhonyy9 (jhonyy9):

???

OpenStudy (anonymous):

Given AB is 9cm but now AB will be 3 ! Thats wer i am confused

jhonyy9 (jhonyy9):

so let be BC = x 3/5 = 9/x so than cros multiplie 3x=45 divide both sides by 3 x= 15 right ?

OpenStudy (anonymous):

Ya!

jhonyy9 (jhonyy9):

yw good luck

OpenStudy (anonymous):

Thanks!! x= 15cm right?

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

Thanks a lot! @jhonyy9 !!

jhonyy9 (jhonyy9):

yw